OpenStudy (anonymous):

Graph the following log equation: y+ log(base1/3)(x-2) Increase or Decreasing? X Intercept? Vertical Asymptote? Domain Range

OpenStudy (anonymous):

\[y = \log_{1/3}(x-2) \]

OpenStudy (anonymous):

Domain: x - 2 > 0 => x > 2

OpenStudy (anonymous):

Range: R

OpenStudy (anonymous):

Log is increasing function

OpenStudy (anonymous):

it has vertical asymptote of x = 0

OpenStudy (anonymous):

What about the x-intercept? @Kanwar245

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

Can you show me how you got that? Thank you so much! @Kanwar245

OpenStudy (anonymous):

x-intercept means where it crosses x-axis, i.e x - 2 = 1 => x = 3 since log 1 = 0

OpenStudy (anonymous):

log (x-2) = 0 when x - 2 = 1

OpenStudy (anonymous):

Thank you so much! @Kanwar245

OpenStudy (anonymous):

How would I sketch the graph? @kanwar245